# Jianwei Li & Jung-Eun Kim

### v2

@@ 5 Broader Impact, Limitations, and Conclusion
Limitation and Future Work. Our investigation into reallocating redundant units for safety purposes is limited to studying the effects of SFT on limited alignment datasets. While this is sufficient to offer indirect support for SSAH, future research should expand this line of inquiry by exploring broader datasets and alternative alignment strategies to further validate and generalize our findings.
@@
Conclusion. This paper distinguishes safety alignment from general alignment and addresses the three key questions: How does safety alignment affect model behavior? Why are safety mechanisms brittle? and How to mitigate the side effects?. By answering these questions, safety alignment is demonstrated as an implicit binary classification task.
Appendix A Appendix: SUPERFICIAL SAFETY ALIGNMENT HYPOTHESIS
@@
A.1 Superficial Alignment Hypothesis in LIMA.
@@
A.2 Model Configuration and Training Details
@@
For the instruction-following ability, we evaluated the helpfulness of the model using the MT-Bench benchmark (Zheng et al., 2023), which assesses the general utility and coherence of model responses across a wide range of tasks. Importantly, we use GPT-4 as a judge to evaluate the helpfulness of the model’s generated responses. Specifically, GPT-4 was used to compare the outputs of our trained models against standard task prompts in MT-Bench and assign scores based on response quality, relevance, and overall helpfulness. To evaluate the safety performance of the models, we employed two complementary benchmarks: Adv-bench and HEx-PHI. These benchmarks were chosen to comprehensively assess the models’ ability to handle malicious or unsafe queries. To save space, please refer to Sec. B.5 for more details about these two datasets and the corresponding metrics.
@@
Appendix B Appendix: Less is More for Safety Alignment
@@
B.1 Definition of Attribute Groups and Categorization Process
@@
Datasets Used for Computing $\mathbf{I_{U}}$ and $\mathbf{I_{S}}$. To identify safety-critical regions in the model, we follow Wei et al. (2024b) to prepare two types of datasets: safety dataset, for attributing safety-related behaviors, and utility dataset, for attributing utility-related behaviors. Each dataset is structured in a (prompt, response) format. Specifically, the safety dataset is compiled using harmful instructions from AdvBench (Zou et al., 2023a). We also divide AdvBench into AdvBencheval (100 instructions for evaluation) and AdvBenchattr (420 instructions for attribution). We prompt Llama2-7B-chat with AdvBenchattr, collecting responses that refrain from following harmful instructions. For the utility dataset, we filter out safety-related (prompt, response) pairs using sensitive phrase matching (Qi et al., 2023) from Alpaca-Cleaned, a refined version of the Alpaca dataset (Taori et al., 2023).
@@
B.2 Evaluating the Impact on Both Utility and Safety Performance
To evaluate the impact of pruning on both utility and safety performance, we measure the model’s performance using established benchmarks for both attributes. Our approach closely follows the methods used by Sun et al. (2023); Wei et al. (2024b); Zou et al. (2023b), with adaptations to focus on the specific aspects of utility and safety in safety alignment.
@@ B.2 Evaluating the Impact on Both Utility and Safety Performance
B.3 Model Pruning Details and Structured Components
To implement structured pruning, we follow the method proposed by Li et al. (2024); An et al. (2024). The pruning process targets specific structured components (neurons or channels) within the depth-2 modules of the transformer architecture, which includes both attention and feedforward layers. A depth-2 module is represented as $f(X)=B\sigma(AX)$, where $A$ and $B$ are weight matrices. This paper focuses on the inner channel pruning (please refer to Fig. 1 in Li et al. (2024)): pruning the input channels of matrix $B$ and the output neurons of matrix $A$. This allows us to directly reduce the number of active channels and neurons in both the feedforward and attention mechanisms, ensuring that less important components (those with low variance) are removed.
@@ B.3 Model Pruning Details and Structured Components
This structured pruning approach, combined with importance score normalization, ensures that the pruned units correspond to meaningful portions of the model, and the results from various pruning ratios provide insight into the essential number of units required for maintaining safety and utility performance. Although our method for identifying safety-critical components shares similarities with the design proposed by Wei et al. (2024b), we utilize a different functional structure. Crucially, our approach has been shown to maintain safety performance under fine-tuning attacks, a result that previous methods were unable to achieve.
B.4 Attribute Transfer During Fine-Tuning
@@
B.5 Experimental Setup for Freezing Safety-Critical Components
@@
    AdvBench (Zou et al., 2023a): AdvBench is a benchmark designed to test a model’s vulnerability to adversarial instructions. It contains prompts specifically crafted to elicit unsafe or harmful outputs. We use two evaluation methods with AdvBench. First, we employ a keyword-matching strategy (the original evaluation method) to detect whether the model produces unsafe outputs. Second, we introduce a new evaluation method using Llama3-Guard Inan et al. (2023), where we treat the Llama3 model as a safety arbiter (or ”judge”) to assess the safety of the outputs. The final evaluation metric is the Attack Success Rate (ASR), with higher ASR indicating a more dangerous model.
@@
B.6 Details of Redundant Units and Alignment Budget

### v3

@@ 5 Conclusion
In conclusion, this paper distinguishes safety alignment from general alignment and addresses the three key questions: How does safety alignment affect model behavior? Why are safety mechanisms brittle? and How to mitigate the side effects?. By answering these questions, safety alignment is demonstrated as an implicit binary classification task.
Appendix A Appendix: More Related Works and Baselines
A.1 Overview of Safety-Unit and Safety-Alignment Work
In this subsection, we compare our work with the following studies: Kim et al. (2025); Zhao et al. (2025); Li et al. (2024b); Chen et al. ; Lu et al. (2025); Yi et al. (2025); Chen et al. (2024); Bhardwaj et al. (2024); Chen et al. (2024); Zou et al. (2024); Hsiung et al. (2025); Liu et al. (2025b; b); Xiao et al. (2025); LLMS ; Huang et al. (2024a); Liu et al. (2025a); Eiras et al. (2024); Huang et al. (2024b); Hsu et al. (2024). For clarity, we group them into three categories: (i) safety neuron/layer identification and manipulation, (ii) data and task-distribution analyses, and (iii) optimization-centric defenses and training recipes.
Table 6: Comparison with representative safety-neuron/layer identification and manipulation methods.
Work Evidence for safety traceability Granularity Identification method Intervention recipe
Ours Attribute-transfer analysis and pruning-based causal tests Neuron & channel level Top-$K$ ranking on $(I_{\text{safety}}-I_{\text{utility}})$ Freeze SCU + $X\%$ CU
Kim et al. (2025) Hyperparameter ablation and causal tests Weight level Exponential moving average over all parameters EMA-based model merging
Zhao et al. (2025) Deactivation causal tests Neuron & channel level Top-$K$ by safety-importance score $I_{\text{safety}}$ Freeze (safety neurons minus foundational neurons)
Li et al. (2024b) Cosine similarity of layer outputs for normal vs. malicious queries Layer/block level Parameter scaling guided by over-rejection phenomenon Freeze safety layers
Chen et al. Activation-patching causal tests Neuron & activation level Activation contrasting + dynamic activation patching Activation patch
Lu et al. (2025) No explicit causal evidence Weight & activation level Safety-bounded weight-delta optimization Weight-delta curation + compensation
Yi et al. (2025) Selective causal restoration Weight level Low-rank projection Neuron patch
Bhardwaj et al. (2024) Weight-delta causal tests Weight level Weight comparison across models Weight patch
Zou et al. (2024) No explicit causal evidence Representation level Circuit-breaking on unsafe representations Circuit breaking
Safety neuron/layer identification and manipulation
Kim et al. (2025) views harmful fine-tuning as an optimization issue and proposes using EMA-based model merging to construct a more stable optimization path that reduces safety degradation. Zhao et al. (2025) finds that fewer than 1% of neurons, mainly in early attention layers, are safety-critical and introduces SN-Tune/RSN-Tune to protect or retrain these safety neurons. Li et al. (2024b) identifies a small contiguous block of “safety layers” that are crucial for distinguishing harmful from benign queries and shows that freezing these layers’ gradients mitigates fine-tuning attacks. Chen et al. ; Chen et al. (2024) locate safety neurons via activation contrast and use dynamic activation patching on roughly 5% of neurons to causally restore most of the model’s original safety performance. Lu et al. (2025) proposes Safe Delta, which selects and adjusts weight deltas to bound safety loss while preserving utility, followed by a safety compensation vector to recover residual safety. Yi et al. (2025) builds a safety reference model to identify safety-critical neurons whose behavior has drifted and selectively restores them via neuron-level patches to realign safety after harmful fine-tuning. Bhardwaj et al. (2024) treats safety as a direction in weight space and restores it by adding a learned “safety vector” (a safety-related weight delta) to a fine-tuned but unsafe model. Zou et al. (2024) introduces “circuit breakers,” which directly intervene on unsafe internal representations (hidden states) to cut off harmful computational pathways before they produce unsafe outputs. We summarize how these methods differ from ours along four axes—evidence for safety traceability, granularity of safety units, identification method, and intervention recipe—in Table 6.
Moreover, our task of interest and problem formulation are conceptually distinct from these safety-neuron/layer works. Most of them are primarily concerned with preventing fine-tuning attacks by identifying safety neurons or layers and proposing concrete protection mechanisms (e.g., freezing, patching, gating) to preserve safety. By contrast, our work focuses on understanding how safety alignment operates within LLMs. We introduce SSAH, which models safety alignment as an implicit safety-related binary decision process (fulfill vs. refuse), and we extend this view to a non-superficial variant for token-level and jailbreak analysis. (However, it may not be used to analyze the backdoor attack (Li and Kim, 2026).) We then provide probing experiments and attribute-transfer analyses that support SSAH and its “less-is-more” implication. Within this framework, the identification of SCUs, UCUs, CUs, and RUs, together with the freeze-SCU/CU, fine-tune-RU intervention, is primarily used as a mechanistic testbed to provide additional indirect evidence for SSAH; they secondarily act as a practical defense against fine-tuning attacks.
Data and task-distribution analyses
Another line of work focuses on alignment data and task distributions; the analysis is typically conducted at the data or feature level (LLMS, ; Liu et al., 2025b). These approaches argue that downstream fine-tuning overlaps safety-relevant features and thus entangles utility and safety (Hsiung et al., 2025). While informative, such explanations are indirect and largely inductive: they infer mechanisms from distributional overlap rather than pinpointing the internal units responsible for safety degradation (Xiao et al., 2025; Eiras et al., 2024). In contrast, SSAH identifies a direct mechanism with explicit empirical evidence that is closer to the model’s computation. By probing internal units and categorizing them as SCU/UCU/CU/RU based on attribute changes before and after SFT or harmful fine-tuning, we reveal a role-reallocation mechanism: during task adaptation, the model often acquires the desired attribute (utility or safety) by converting computing units that originally contributed to the other attribute. This unit-level migration (e.g., CU $\rightarrow$ UCU, SCU $\rightarrow$ RU) provides structural and causal evidence for brittleness and alignment tax, offering a direct internal account that complements these data-level perspectives.
Optimization-centric defenses and training recipes
A third line of work proposes optimization-centric defenses or training recipes (Huang et al., 2024a). For example, Booster and Targeted Vaccine adopt adversarial-training-style perturbations (global or layer-targeted) to harden models against downstream safety drift (Liu et al., 2025a)Huang et al. (2024b) formulates safety preservation as a bi-state optimization problem with a stability term; Safe LoRA constrains adapters by projecting toward a safety-aligned subspace to reduce interference (Hsu et al., 2024). These approaches offer practically useful mechanisms but do not explicitly explain and verify why safety collapses under benign finetuning. Within SSAH, safety is described as an implicit binary classification task realized by SCUs and CUs, and fine-tuning degrades safety or utility through attribute reallocation among these units. Consequently, the freeze SCU/CU, fine-tune RU rule follows mechanistically from the framework, rather than being a standalone heuristic, and can also guide where to perturb, regularize, or project when adopting optimization-based defenses (Huang et al., 2024a; Liu et al., 2025a).
In summary, across these three categories of prior work, SSAH provides a foundational hypothesis and framework that (i) explains how safety alignment reshapes model behavior, (ii) verifies and interprets the observed less-is-more effect and safety brittleness via unit-role dynamics, and (iii) yields a minimal, principled intervention—freezing SCU/CU and fine-tuning RU—that complements existing neuron/layer-, data-, and optimization-based methods.
A.2 Empirical Comparisons with Safety-Unit Freezing Baselines
We further compare our method with two closely related baselines, Zhao et al. (2025) and Li et al. (2024b), which identify safety-relevant components and freeze them to mitigate fine-tuning attacks. We choose these two because all three approaches share the same high-level recipe—identify safety units and freeze them—but differ in (1) identification criteria and (2) granularity of the safety units.
Zhao et al. (2025) operates at the neuron/channel level. They use a safety-related dataset to detect (within each layer) the neurons or channels whose deactivation causes the largest change in the norm of layer output. These identified components are treated as “safety neurons”. The same procedure is then applied to a Wikipedia corpus to detect “foundation neurons”. During fine-tuning, they freeze the subset of safety neurons that do not overlap with foundation neurons. In contrast, Li et al. (2024b) operates at the layer/block level. It scales the hidden-state signals for candidate block ranges and evaluates over-rejection on a synthetic dataset to identify a contiguous segment of “safety layers” whose freezing best preserves safety. For LLaMA-2-7B-Chat, Li et al. (2024b) recommends freezing blocks [6–14], which together correspond to about 28.1% of all parameters. Because these two studies work at very different granularities, it would be unfair to force them into a single unified comparison setting; instead, we design two separate comparison setups.
In the first setting, we compare Zhao et al. (2025) with our method at the neuron level. We perform benign fine-tuning on Alpaca and GSM8K and then evaluate safety performance on AdvBench and HEx-PHI before and after this adaptation. To ensure fairness, (1) for the safety-related dataset used to identify safety neurons in Zhao et al. (2025), we adopt the same safety dataset as in our method; (2) for foundation neurons in Zhao et al. (2025), we follow their design and use a Wikipedia corpus (rather than our instruction-tuned utility dataset); (3) for the freezing ratio in Zhao et al. (2025), we follow their 0.4% selection rule on LLaMA-2-7B-Chat; for our method, we keep the default setting in the main paper, freezing all SCUs at 1.3% plus the top 6% of CUs on LLaMA-2-7B-Chat; and (4) we additionally include a scaled-up variant of Zhao et al. (2025), where we enlarge their safety-neuron selection to 6.3% to match our freezing budget more closely.
Importantly, Zhao et al. (2025) selects safety neurons solely by safety importance $I_{\text{safety}}$. It defines safety neurons as the minimal number of neurons whose deactivation suffices to noticeably degrade safety while not harming it beyond a threshold. This yields a minimal set under that constraint. However, the minimum set of neurons required to destroy safety by deactivation is often insufficient to implement robust safety guardrails. Consequently, when they further remove overlap with foundation neurons in the freezing stage, the remaining safety neurons are insufficient to prevent safety degradation (Their strategy will lead to an inflexible and often trivial freezing ratio, they also note that “a complete harmful score reduction to 0.0 is not achievable due to an insufficient number of non-overlapping safety neurons”). In contrast, we define SCUs as the maximal set of units whose removal degrades safety while not degrading utility. In practice, we rank units by the relative importance score $I_{\text{safety}}-I_{\text{utility}}$ and choose the largest subset that satisfies safety $\downarrow$ and utility $\approx$ steady. This maximal view (i) directly separates SCU from CU and (ii) provides a flexible way to choose the freezing ratio, as we can gradually expand the budget according to the score $I_{\text{safety}}-I_{\text{utility}}$. The quantitative results for this setting are reported in the Tab. 78 for Alpaca and GSM8K.
Table 7: Safety performance on Alpaca fine-tuning: comparison between our method and work Zhao et al. (2025).
Bench Judge Initial Full Finetuned Ours Zhao et al. (2025) Zhao et al. (2025) (6.3%)
Adv keyword 0.19% 5.30% (+5.11%) 2.96% (+2.77%) 5.19% (+5.00%) 4.42% (+4.23%)
Adv llama3-guard 0.19% 2.69% (+2.50%) 1.65% (+1.46%) 4.23% (+4.04%) 3.46% (+3.27%)
HEx-PHI GPT-4 (1–5) 1.05 1.79 (+0.74) 1.39 (+0.34) 1.72 (+0.67) 1.54 (+0.49)
HEx-PHI llama3-guard 2.42% 18.40% (+15.98%) 12.12% (+9.70%) 13.33% (+10.91%) 12.42% (+10.00%)
Table 8: Safety performance on GSM8K fine-tuning: comparison between our method and work Zhao et al. (2025).
Bench Judge Initial Full Finetuned Ours Zhao et al. (2025) Zhao et al. (2025) (6.3%)
Adv keyword 0.19% 5.38% (+5.19%) 1.92% (+1.73%) 3.85% (+3.66%) 2.88% (+2.68%)
Adv llama3-guard 0.19% 5.77% (+5.58%) 1.35% (+1.16%) 3.65% (+3.46%) 2.50% (+2.31%)
HEx-PHI GPT-4 (1–5) 1.05 1.60 (+0.50) 1.37 (+0.32) 1.48 (+0.43) 1.44 (+0.39)
HEx-PHI llama3-guard 2.42% 17.88% (+15.46%) 11.52% (+9.10%) 15.75% (+13.33%) 12.42% (+10.00%)
In the second setting, we compare Li et al. (2024b) with our method at a matching parameter budget at the layer/block level. Again, we perform benign fine-tuning on Alpaca and GSM8K, and then evaluate safety on AdvBench and HEx-PHI before and after adaptation. Here, (1) for Li et al. (2024b), we follow their recommendation and freeze blocks 6–14 on LLaMA-2-7B-Chat, which corresponds to approximately 28.1% of all parameters; and (2) for our method, to ensure a fair parameter budget, we increase the CU freezing ratio from 6% to 26.8%, so that SCU plus CU together also freeze about 28.1% of parameters. The results for this setting are in Tab 9 & 10.
Table 9: Safety performance on Alpaca fine-tuning: comparison between our method and work Li et al. (2024b) under a matched parameter budget (28.1% frozen).
Bench Judge Initial Full Finetuned Ours Li et al. (2024b)
Adv keyword 0.19% 5.30% (+5.11%) 2.30% (+2.11%) 3.85% (+3.66%)
Adv llama3-guard 0.19% 2.69% (+2.50%) 1.15% (+0.96%) 3.46% (+3.27%)
HEx-PHI GPT-4 (1–5) 1.05 1.79 (+0.74) 1.26 (+0.21) 1.41 (+0.36)
HEx-PHI llama3-guard 2.42% 18.40% (+15.98%) 8.18% (+5.76%) 12.12% (+9.70%)
Table 10: Safety performance on GSM8K fine-tuning: comparison between our method and work Li et al. (2024b) under a matched parameter budget (28.1% frozen).
Bench Judge Initial Full Finetuned Ours Li et al. (2024b)
Adv keyword 0.19% 5.38% (+5.19%) 1.73% (+1.54%) 2.50% (+2.31%)
Adv llama3-guard 0.19% 5.77% (+5.58%) 1.15% (+0.96%) 2.30% (+2.11%)
HEx-PHI GPT-4 (1–5) 1.05 1.60 (+0.50) 1.31 (+0.26) 1.43 (+0.38)
HEx-PHI llama3-guard 2.42% 17.88% (+15.46%) 10.30% (+7.88%) 11.82% (+9.40%)
Overall, under comparable or even stricter parameter budgets, our freezing strategy yields lower attack success rates than the neuron-level baseline Zhao et al. (2025) and the layer-level baseline Li et al. (2024b). This advantage arises from two main factors: (1) an identification principle that selects SCUs via a maximal-margin criterion on $I_{\text{safety}}-I_{\text{utility}}$, in contrast to the minimal knockout sets in Zhao et al. (2025), which are incomplete and utility-agnostic; and (2) a finer granularity and more direct targeting of safety–utility attributions than Li et al. (2024b), which operates at the coarse block level and is guided by an over-rejection proxy that is indirectly tied to safety metrics.
Appendix B Appendix: SUPERFICIAL SAFETY ALIGNMENT HYPOTHESIS
@@
B.1 Superficial Alignment Hypothesis in LIMA.
@@
B.2 Model Configuration and Training Details
@@
For the instruction-following ability, we evaluated the helpfulness of the model using the MT-Bench benchmark (Zheng et al., 2023), which assesses the general utility and coherence of model responses across a wide range of tasks. Importantly, we use GPT-4 as a judge to evaluate the helpfulness of the model’s generated responses. Specifically, GPT-4 was used to compare the outputs of our trained models against standard task prompts in MT-Bench and assign scores based on response quality, relevance, and overall helpfulness. To evaluate the safety performance of the models, we employed two complementary benchmarks: Adv-bench and HEx-PHI. These benchmarks were chosen to comprehensively assess the models’ ability to handle malicious or unsafe queries. To save space, please refer to Sec. C.5 for more details about these two datasets and the corresponding metrics.
@@
Appendix C Appendix: Less is More for Safety Alignment
@@
C.1 Definition of Attribute Groups and Categorization Process
@@
Datasets Used for Computing $\mathbf{I_{U}}$ and $\mathbf{I_{S}}$. To identify safety-critical regions in the model, we follow Wei et al. (2024) to prepare two types of datasets: safety dataset, for attributing safety-related behaviors, and utility dataset, for attributing utility-related behaviors. Each dataset is structured in a (prompt, response) format. Specifically, the safety dataset is compiled using harmful instructions from AdvBench (Zou et al., 2023a). We also divide AdvBench into AdvBencheval (100 instructions for evaluation) and AdvBenchattr (420 instructions for attribution). We prompt Llama2-7B-chat with AdvBenchattr, collecting responses that refrain from following harmful instructions. For the utility dataset, we filter out safety-related (prompt, response) pairs using sensitive phrase matching (Qi et al., 2023) from Alpaca-Cleaned, a refined version of the Alpaca dataset (Taori et al., 2023).
@@
C.2 Evaluating the Impact on Both Utility and Safety Performance
To evaluate the impact of pruning on both utility and safety performance, we measure the model’s performance using established benchmarks for both attributes. Our approach closely follows the methods used by Sun et al. (2023); Wei et al. (2024); Zou et al. (2023b), with adaptations to focus on the specific aspects of utility and safety in safety alignment.
@@ C.2 Evaluating the Impact on Both Utility and Safety Performance
C.3 Model Pruning Details and Structured Components
To implement structured pruning, we follow the method proposed by Li et al. (2024a); An et al. (2024). The pruning process targets specific structured components (neurons or channels) within the depth-2 modules of the transformer architecture, which includes both attention and feedforward layers. A depth-2 module is represented as $f(X)=B\sigma(AX)$, where $A$ and $B$ are weight matrices. This paper focuses on the inner channel pruning (please refer to Fig. 1 in Li et al. (2024a)): pruning the input channels of matrix $B$ and the output neurons of matrix $A$. This allows us to directly reduce the number of active channels and neurons in both the feedforward and attention mechanisms, ensuring that less important components (those with low variance) are removed.
@@ C.3 Model Pruning Details and Structured Components
This structured pruning approach, combined with importance score normalization, ensures that the pruned units correspond to meaningful portions of the model, and the results from various pruning ratios provide insight into the essential number of units required for maintaining safety and utility performance. Although our method for identifying safety-critical components shares similarities with the design proposed by Wei et al. (2024), we utilize a different functional structure. Crucially, our approach has been shown to maintain safety performance under fine-tuning attacks, a result that previous methods were unable to achieve.
C.4 Attribute Transfer During Fine-Tuning
@@
C.5 Experimental Setup for Freezing Safety-Critical Components
@@
    AdvBench (Zou et al., 2023a): AdvBench is a benchmark designed to test a model’s vulnerability to adversarial instructions. It contains prompts specifically crafted to elicit unsafe or harmful outputs. We use two evaluation methods with AdvBench. First, we employ a keyword-matching strategy (the original evaluation method) to detect whether the model produces unsafe outputs. Second, we introduce a new evaluation method using Llama3-Guard Inan et al. (2023), where we treat the Llama3 model as a safety arbiter (or "judge") to assess the safety of the outputs. The final evaluation metric is the Attack Success Rate (ASR), with higher ASR indicating a more dangerous model.
@@
C.6 Details of Redundant Units and Alignment Budget